Minute of KAGRA commissioning meeting
- Date: 2018/7/17 11AM~12PM
- Attendance: Miyakawa, Miyamoto, Aso, Nakano, Yokozawa
- eZuce: Shoda, Izumi, Ushiba
Clipping loss (Izumi)
- 20mm misalignment is acceptable for beam cetering.
- 40mm misalignment is acceptable for WAB cetering.
- Could be factor of 2 different in loss.
- Side of screw of frame was almost touching in the room temperature and after starting cooling it was actually touching.
- Then it was found that F2 was almost dead. Even VIS team pulled wire but LVDT did not work.
- Then pump down started without F2 fix.
- New measurement shows it seems to be OK except for 2.4Hz peak probably due to the dead of F2.
- Cooling started the last Friday.
- Heat link does not affect so much.
- Fiber was broken just before moving the laser to Kamioka.
- We have 1 + 1spare = 2.
- Low power below 15A has many peaks.
- Low frequency RIN is proportional to the power.
- Movement to Kamioka will be the 24th of July.
- PMC and iMC layout including MZ is being changed.
- Cable cleaning on BS rack is being continued.
- pile of AA or HPCD can be hot.
- Next meeting: 7/24.